# Shape-specific fluctuations of an active colloidal interface

### v1

@@ IV Interface fluctuations
Figure 2: Schematic of dynamical evolution of colloidal chain. From left to right: evolution of the chain at different times $t=0$, until the final $t=T$. At times above $\tau$ and $\tau_{f}$ the chain acquires the steady-state C-shape. Note that the chain propels in the positive $x$ direction, and that the time points are not linearly scaled. The inset on each show the monomer orientations for a particular segment at each time. $N^{\prime}$is labelled in red (middle panel), along with $\Delta h$ (rightmost panel) and $\theta$ (in inset).
For the C-shape chain propelling in the positive x-direction, one can define a height function $h(y)$ along the chain. An example of its temporal evolution is shown in Fig. 2. To differentiate the height of this shape from other typically measured shapes (circular and flat), let us call this the “C-interface” (CI). One can thus choose a particular length segment along the chain and compute the height fluctuations, and measure its root-mean-squared
@@
for the CI. Here, $\langle\rangle$ is taken over both the segment and realizations. We can do the same for the angular fluctuations, and define
@@
For the orientational sector, we find that, though there is a growth followed by a plateau, there is a persistent smoothening phenomena, where there is no system-size dependent saturation time scale, but instead a system-size dependent scaling of the fluctuations at all times, which is instead negative (hence smoothness). We discuss this in detail below. In this case, the scaling reads:
@@
with the important difference between the two being the $N$ dependence of the characteristic saturation time of fluctuations $t^{*}$, as explained below. Given (15), $W_{\theta}^{\text{SS}}$ can be equivalently be taken to be the time average over the entire simulation; we denote this as $\langle W_{\theta}\rangle_{t}$ wherever used.
We note that the results in this Section are based on the assumption of instantaneous chemical depositions and the initial conditions of Eq. (4); we display in Appendix .7 that similar growth exponents are obtained when chemical trails are present; where the C-shape is instead universal (independent of initial conditions) [28].
IV.1 Specific shape about which to measure fluctuations
We emphasize here that the fluctuations reported in this work correspond to those exclusively about the C-shape. In the “stiff” phase $\Lambda\rightarrow\infty$ these would trivially reproduce an early-time growth for $W_{h}$, similar to that for a driven EW model, as shown explicitly in Appendix .1. Fluctuations could, in principle, be computed about various other DSS as well. To emphasize the uniqueness of this shape let us list some generic properties of a DSS:
  1. 1.
    Continuously varying orientations along the chain
  2. 2.
    Finite curvature along chain (or the averaged local curvature)
  3. 3.
    Ballistically propelling in one direction
  4. 4.
    Positional and orientational symmetry along the body(y) axis
As explained in Appendix .4, there are various other DSS that meet some of the above requirements but not all. The aforementioned “stiff” DSS satisfies only (i), (iii) and (iv). Another example of this is that of a C with a “pinch” defect, where defects in the orientation along the chain renders a finite number of segments of the chain to exhibit a C shape (Fig. 7(D) in Appendix .4). Here $\tau_{f}\sim\tau_{t}\sim\tau$, thus sharp reorientations occur alongside a positional relaxation, allowing for occasional orientational defects, all of this whilst global curvature sets in. Indeed, the C-shape is unique in that it is the only DSS that meets requirements (i)-(iv), with its propulsion direction to the body axis. It is this DSS about which the fluctuations are computed.
@@ IV.1 Specific shape about which to measure fluctuations
To compute the quantities $W_{h}(t)$ and $W_{\theta}(t)$ [Eq. (8) and (9)] respectively, we discuss two possibilities. As displayed in Fig. 2, there are two separate segments about which fluctuations are measurable. Let us label the length of this segment to be $N^{\prime}$. The first obvious segment corresponds to $N^{\prime}=N$, corresponding to the average over the entire CI. The second corresponds to that of a locally flat region; here $N^{\prime}<<N$, which is taken to be sufficiently far from the edges of the chain. Though established results exist for fluctuations about curved segments of an interface [9], these typically correspond to that of a (subset of a) circular interface, thus different from that of our CI (see Appendix .3). For the latter, it is to be noted that fluctuations of growing (circular) interfaces are typically measured in this regime [31, 32] ; the scaling properties in this region are almost always equal to that of globally measured quantities - for instance in [16] for a flat surface and [31] for a curved surface, and indeed we extend this equivalence to the CI.
@@
Figure 3: Scaling of height fluctuations. A $W_{h}$ for different chain lengths NN, with the three exponents $\beta_{h}$ indicated via dashed black line. The same scaling is plotted in the inset, with $t_{h}^{*}(N)$ labelled in dashed vertical line. B Roughness scaling for $W_{h}^{\text{SS}}(N)$. Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
IV.2 Height fluctuations
@@
(i) Early-time regime I. This corresponds to times $t<<\tau$ (first few time steps of simulation). Here the growth of the interface height is dominated purely by noise, the mean-squared displacement of the height scales as $t$, and thus $\beta^{(0)}_{h}\sim 0.5$.
(ii) Early-time regime II. Now the colloids start to feel interactions from their respective neighbors which leads to slowing down of the pure diffusive growth. Deterministic chemical self-interactions, though, are yet to set in. Here, $\beta^{(1)}_{h}\approx 0.25$. These first two regimes correspond to those observed in the EW model [33]. For $t<\tau_{f}$, it is straightforward to show that our chain approximates an EW interface, as shown in Appendix .1. These early time scalings correspond to (as expected) a purely diffusive regime, equivalent to random deposition of particles on a substrate accompanied by surface diffusion. An alternate way to characterize the early time growth is via the probability distribution of $W^{2}_{h}$; this (and a comparison with the corresponding EW distribution) is presented in Appendix .1.
(iii) Super-ballistic regime: In this regime $t>\tau_{f}$ with $t<t_{h}^{*}$. During this time, the self-phoretic interactions cause the chain to morph into a C-shape topology. As a result, $W_{h}$ swells up to its steady-state value, with a super-ballistic exponent (quoted below). Let us study the behaviour of these fluctuations by defining a correlation length $l_{c}$. Via dimensional analysis, $l_{c}\sim b(\frac{t}{\tau})^{\frac{1}{z_{h}}}$. In this region $l_{c}$ typically corresponds to a finite fraction along the interface. For instance, at $t=10\tau$ time steps into the simulation (at the onset of swelling, see inset of Fig. 3(A)), $l_{c}\approx 53b$. These correlations further build-up until the next regime is reached.
(iv) Steady-state regime. This regime corresponds to $t>>\tau_{f}$ and $t>>t_{h}^{*}$. We identify $t_{h}^{*}=N^{z_{h}}\tau$, a system-size dependent characteristic timescale over which the fluctuations reach a steady-state (see inset of Fig. 3)(A). For example for $N=256$ (Fig.3(A) inset, green), $t_{h}^{*}\approx 27.6\tau\sim 10^{2}\tau_{f}$. Here, the fluctuations cease to grow and $W_{h}$ reaches a plateau. The aforementioned correlations now scale as $(\frac{t}{\tau})^{\frac{1}{z_{h}}}\sim N$ (spanning the entire chain length).
@@
From (18) and Fig. 3, we find that regimes (iii) and (iv) exhibit a clear FV scaling with a roughness exponent $\alpha_{h}$ and a dynamic exponent $z_{h}=\frac{\alpha_{h}}{\beta_{h}}\approx 0.6$. To our knowledge this set of exponents does not correspond to any previously reported universality class. We thus interpret this as a novel “CI roughness” scaling, distinct to those found for either circular or flat interfaces [31, 15]. The reason for the super-ballistic exponent arises from the fact that as the C-shape formation takes place, displacements of each monomer from the mean is enhanced in either direction (along the x-axis) along the chain. These squared deviations from the mean height are in addition symmetric along the y-axis. $l_{c}$ is clearly enhanced due to this symmetry. It is to be noted that in a conventional ballistic scaling one simply has uni-directional deviations along the propulsion axis; in our system this is applicable to the mean $h(y)$; clearly $l_{c}$ in this case is smaller than for a symmetric C-shape. We suggest that these combined effects give rise to the super-ballistic growth exponent $\beta_{h}$. The presence of a super-ballistic exponent can be rationalized via a mean-field analysis of the EOM 2. This is presented in Appendix .2. It is to be noted that varying $\tau_{r}$ changes the time at which super-ballistic growth sets in (exponent and roughness is unchanged).
The roughness exponent $\alpha_{h}$, on the other hand, indicates the relative growth of height fluctuations when $l_{c}$ spans the system size. Here, the system is is at $\sim 10^{2}\tau_{f}$ and the C-shape is fully formed. The value that we report is thus quite appreciably larger than those reported in EW [33], KPZ [34] and most numerical deposition models such as for instance the solid-on-solid growth model [16].
IV.3 Orientational fluctuations
@@
Figure 4: Scaling of the orientation fluctuations. A $W_{\theta}$ for different chain lengths NN, with the dynamic $\beta_{\theta}$ labelled. The same scaling is plotted in the inset, with $t^{*}$ labelled in dashed vertical line. B (Anti-) Roughness scaling for $W_{\theta}^{\text{SS}}(N)$. C Orientation profile of the C-shape, taken at different time points. D Dynamical evolution of $\theta_{i}$ for selected monomers ii. Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$, and $N=311$ (for panels C and D).
@@
The ballistic regime arises due to the purely deterministic contribution to the orientational changes that sets in during $t\sim 5\tau_{f}\approx\tau$. The emergence of the ballistic regime can be intuitively appreciated by observing the orientation profile $\theta_{i}$ in Fig 4(C) for $N=312$. We see that at all times $\theta_{i}$ maintains strong correlations across the chain, with $\theta_{i}=-\theta_{N-i}$ (anti-symmetry; here $i$ is the monomer index). Such an anti-symmetric profile of the orientation field along the y-axis, along with the dynamics thus enforces $l_{c}$ to span the entire chain length. Further, as the steady-state is approached, a large number of orientations remain close to their initial value (4(d)). For instance, if one takes a threshold $\theta^{*}=0.2$ (corresponding to $|\frac{\max(\theta)-\theta^{*}}{\max(\theta)}|\approx 0.92$ deviation from the edge monomer angles), we find that a fraction of $P(\theta_{i}<\theta^{*})\approx.89$ of monomers on the chain remain roughly equal to their initial value, with the remaining $\approx 11\%$ drastically varying at the edges. This notion of orientational “stiffness” is also discussed in [24]. We note that here $\tau_{r}\rightarrow\infty$; the results for finite $\tau_{r}$ are presented in Appendix .6; there is in addition an early time diffusive regime, though that does not affect the results presented here  [^1] As explained in [28], sufficiently large rotational noise prohibits the C-shape formation. Thus, “positive $D_{r}$” wherever mentioned here implicitly assumes $\frac{\tau_{f}}{\tau_{r}}=\frac{b^{3}D_{r}}{\chi_{r}}<<1$. Here, $\frac{\tau_{f}}{\tau_{r}}=0.01$..
@@
We further report that $W_{\theta}^{SS}$ scales negatively with the chain length. We may thus call this a smoothness exponent. This form of orientational stabilization with the system size, as opposed to conventional “roughness” seen in other growing interfaces (e.g. $\alpha^{\text{EW}}\approx 0.5$ and $\alpha^{\text{KPZ}}\approx 1/3$). The source of emergence for such a smoothening exponent can be appreciating the increased orientational rigidity along the chain, which increases with chain length [24]. The angular distribution along the CI becomes increasingly uniform, thereby making the chain stiffer and more polarized along the x-axis. The RMS deviations from the the mean orientation thus decreases with system size (a larger fraction of the chain is polarized along the x-axis). This is in contrast to the positional steady-state fluctuations, where the mean $h(y)$ varies with chain length, and fluctuations about this exhibits roughness. Note that this is although the chain-averaged curvature vanishes in the thermodynamic limit .3.

### v2

@@ IV Interface fluctuations
For the C-shape chain propelling in the positive x-direction, one can define a height function $h(y)$ along the chain. An example of its temporal evolution is shown in Fig. 1(d). One can thus choose a particular length segment along the chain and compute the height fluctuations, and measure its root-mean-squared
@@
for the C-shape. Here, $\langle\rangle$ is taken over both the segment and realizations. We can do the same for the angular fluctuations, and define
@@
For the orientational sector, we find that, though there is a growth followed by a plateau, there is a persistent smoothening phenomena, where there is no system-size dependent saturation time scale, but instead a system-size dependent scaling of the fluctuations at all times, which is instead negative (hence smoothness). This is further elaborated on below. In this case, the scaling reads:
@@
with the important difference between the two being the $N$ dependence of the characteristic saturation time of fluctuations $t^{*}$, as explained below. We note that given (15), $W_{\theta}^{\text{SS}}$ can be equivalently be taken to be the time average over the entire simulation.
@@
To compute the quantities $W_{h}(t)$ and $W_{\theta}(t)$ [Eq. (8) and (9)] respectively, we discuss two possibilities. There are two separate segments about which fluctuations are measurable. Let us label the length of this segment to be $N^{\prime}$. The first obvious segment corresponds to $N^{\prime}=N$, corresponding to the average over the entire C-shape. The second corresponds to that of a locally flat region; here $N^{\prime}<<N$, which is taken to be sufficiently far from the edges of the chain. Though established results exist for fluctuations about curved segments of an interface [9], these typically correspond to that of a (subset of a) circular interface, thus different from that of the C-shape (see Appendix .5). For the latter, it is to be noted that fluctuations of growing (circular) interfaces are typically measured in this regime [36, 37] ; the scaling properties in this region are almost always equal to that of globally measured quantities - for instance in [16] for a flat surface and [36] for a curved surface. We discuss the specific instances of equivalences and/or differences for the C-shape below.
All the results presented in this paper are computed for $N$ values of $N\in[32,64,100,128,192,200,256,312]$, with curves for specific $N$ displayed (and labelled) in each respective figure. Results for the incorporation of rotational noise are computed for $N\in[64,128,192,256,312]$. Choices for $N^{\prime}$ are stated explicitly in the respective section.
@@
Figure 2: Scaling of height fluctuations. A $W_{h}$ for different chain lengths NN, with the three exponents $\beta_{h}$ indicated via dashed black line, fit within $(t/\tau N^{z_{h}})\in[1,20]$. The same scaling is plotted in the inset, with $t_{h}^{*}(N)$ labelled in dashed vertical line. B Roughness scaling for $W_{h}^{\text{SS}}(N)$, with effective exponent in inset. Note that for the effective exponents $N\in[64,312]$ (excluding the smallest $N=32$, due to forward-difference) Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$.
V Results
A summary of the results in this paper is presented in Table 1, wherein the respective Figures are referenced. It is to be noted that the value of $\beta$ for each respective figure is obtained by obtaining the best collapse of the scaling relations 12 and 14 (by eye). The values for $\alpha$ on the other hand are obtained by a least-squares fit on the steady state values of $W$ for each system size. For each plot of $W$, the fitting windows used for the in-figure $\beta$ line of best fit is mentioned in the caption.
Figure 3: Scaling of the orientation fluctuations. A $W_{\theta}$ for different chain lengths NN, with the dynamic $\beta_{\theta}$ labelled (fit between $(t/\tau N^{z_{h}})\in[10,200]$). The same scaling is plotted in the inset, with $t^{*}$ labelled in dashed vertical line. B (Anti-) Roughness scaling for $W_{\theta}^{\text{SS}}(N)$, with effective exponent in the inset. C Orientation profile of the C-shape, taken at different time points. D Dynamical evolution of $\theta_{i}$ for selected monomers ii. Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$, and $N=311$ (for panels C and D).
V.1 Height fluctuations
@@
  1. (i)
    Early-time regime I. This corresponds to times $t<<\tau$ (first few time steps of simulation). Here the growth of the interface height is dominated purely by noise, the mean-squared displacement of the height scales as $t$, and thus $\beta^{(0)}_{h}\sim 0.5$.
  2. (ii)
    Early-time regime II. Now the colloids start to feel interactions from their respective neighbors which leads to slowing down of the pure diffusive growth. Deterministic chemical self-interactions, though, are yet to set in. Here, $\beta^{(1)}_{h}\approx 0.25$. These first two regimes correspond to those observed in the EW model [34]. For $t<\tau_{f}$, it is straightforward to show that our chain approximates an EW interface, as shown in Appendix .2. These early time scalings correspond to (as expected) a purely diffusive regime, equivalent to random deposition of particles on a substrate accompanied by surface diffusion. An alternate way to characterize the early time growth is via the probability distribution of $W^{2}_{h}$; this (and a comparison with the corresponding EW distribution) is presented in Appendix .2.
  3. (iii)
    Super-ballistic regime: In this regime $t>\tau_{f}$ with $t<t_{h}^{*}$. During this time, the self-phoretic interactions cause the chain to morph into a C-shape topology. As a result, $W_{h}$ swells up to its steady-state value, with a super-ballistic exponent ($\beta_{h}^{(2)}$, quoted below). Let us study the behaviour of these fluctuations by defining a correlation length $l_{c}$. Via dimensional analysis, $l_{c}\sim b(\frac{t}{\tau})^{\frac{1}{z_{h}}}$. In this region $l_{c}$ typically corresponds to a finite fraction along the interface. For instance, at $t=10\tau$ time steps into the simulation (at the onset of swelling, see inset of Fig. 2(A)), $l_{c}\approx 53b$. These correlations further build-up until the next regime is reached.
  4. (iv)
    Steady-state regime. This regime corresponds to $t>>\tau_{f}$ and $t>>t_{h}^{*}$. We identify $t_{h}^{*}=N^{z_{h}}\tau$, a system-size dependent characteristic timescale over which the fluctuations reach a steady-state (see inset of Fig. 2)(A). For example for $N=256$ (Fig.2(A) inset, green), $t_{h}^{*}\approx 27.6\tau\sim 10^{2}\tau_{f}$. Here, the fluctuations cease to grow and $W_{h}$ reaches a plateau. The aforementioned correlations now scale as $(\frac{t}{\tau})^{\frac{1}{z_{h}}}\sim N$ (spanning the entire chain length). System size scaling. Fig. 2(B) shows the scaling of $W_{h}^{SS}$ with the chain length $N$; a roughness exponent of $\alpha_{h}\approx 0.9$ is found. The errorbars on coordinate points are computed by taking the standard deviation of $W_{h}$ over 50 realizations in the steady-state. In the inset the effective exponent $\alpha_{h}^{\text{eff}}=\frac{d\log W_{h}^{SS}}{d\log N}$ is plotted. A moderate convergence is observed to the best fitted exponent of $0.9$ (black dotted horizontal line in inset).
@@
From (18) and Fig. 2, we find that regimes (iii) and (iv) exhibit a clear FV scaling with a roughness exponent $\alpha_{h}$ and a dynamic exponent $z_{h}=\frac{\alpha_{h}}{\beta_{h}}\approx 0.52$. To the best of our knowledge, this set of exponents does not correspond to any previously reported universality class. We thus interpret this as a novel “C-shape roughness” scaling, distinct to those found for either circular or flat interfaces [36, 15]. The super-ballistic exponent can be appreciated by noting that as the C-shape formation takes place, displacements of each monomer from the mean is enhanced in either direction (along the x-axis) along the chain. These squared deviations from the mean height are in addition symmetric along the y-axis. $l_{c}$ is clearly enhanced due to this symmetry. It is to be noted that in a conventional ballistic scaling one simply has uni-directional deviations along the propulsion axis; in our system this is applicable to the mean $h(y)$; clearly $l_{c}$ in this case is smaller than for a symmetric C-shape. We suggest that these combined effects give rise to the super-ballistic growth exponent $\beta_{h}$. The presence of a super-ballistic exponent can be rationalized via a linearised analysis of the EOM 2. This is presented in Appendix .3. It is to be noted that varying $\tau_{r}$ changes the time at which super-ballistic growth sets in (exponent and roughness is unchanged). The roughness exponent $\alpha_{h}$, on the other hand, indicates the relative growth of height fluctuations when $l_{c}$ spans the system size. Here, the system is is at $\sim 10^{2}\tau_{f}$ and the C-shape is fully formed. We note that the value we report is in between the standard roughness encountered in dynamics without conservation [34, 35] and smaller than models with mass conservation [38, 17, 39]. In addition larger than most numerical deposition models such as for instance the solid-on-solid growth model [16]. We discuss these discrepancies further in the Discussion section.
Figure 4: Height and orientational fluctuations for finite $\tau_{r}$. A $W_{h}$ for different chain lengths NN, with the dynamic $\beta_{h}$ labelled (fitting window is $t/(\tau N^{z_{h}})\in[0.5,5]$). Scaling for $W_{h}^{\text{SS}}(N)$, with effective exponent in the inset. C $W_{\theta}$ for different chain lengths NN, with the dynamic $\beta_{\theta}$ labelled (fitting window is $t/(\tau N^{z_{\theta}})\in[4,80]$). D (Anti-) Roughness scaling for $W_{\theta}^{\text{SS}}(N)$, with effective exponent in the inset. Here, $\tau_{f}=0.2$, $\Lambda=0.2$, $\tilde{\Lambda}=0.01$ and $\tau_{r}=70$.
V.2 Orientational fluctuations
@@
The ballistic regime arises due to the purely deterministic contribution to the orientational changes that sets in during $t\sim 5\tau_{f}\approx\tau$. The emergence of the ballistic regime can be appreciated by consideration of the orientation profile $\theta_{i}$ in Fig 3(C) for $N=312$. We see that at all times $\theta_{i}$ maintains strong correlations across the chain, with $\theta_{i}=-\theta_{N-i}$ (anti-symmetry; here $i$ is the monomer index). Such an anti-symmetric profile of the orientation field along the y-axis, along with the dynamics thus enforces $l_{c}$ to span the entire chain length. Further, as the steady-state is approached, a large number of orientations remain close to their initial value (3(d)). For instance, if one takes a threshold $\theta^{*}=0.2$ (corresponding to $|\frac{\max(\theta)-\theta^{*}}{\max(\theta)}|\approx 0.92$ deviation from the edge monomer angles), we find that a fraction of $\approx.89$ of monomers on the chain remain roughly equal to their initial value, with the remaining $\approx.11$ drastically varying at the edges (“orientational stiffness”[25]). We note that here $\tau_{r}\rightarrow\infty$; the results for finite $\tau_{r}$ are presented in Appendix V.3; there is in addition an early time diffusive regime, though that does not affect the results presented here  [^1] As explained in [29], sufficiently large rotational noise prohibits the C-shape formation. Thus, “positive $D_{r}$” wherever mentioned here implicitly assumes $\frac{\tau_{f}}{\tau_{r}}=\frac{b^{3}D_{r}}{\chi_{r}}<<1$. Here, $\frac{\tau_{f}}{\tau_{r}}=0.01$..
@@
We further report that $W_{\theta}^{SS}$ scales negatively with the chain length. We may thus call this a smoothness exponent. This form of orientational stabilization with the system size, as opposed to conventional “roughness” seen in other growing interfaces. The source of emergence for such a smoothening exponent is thus the increased orientational rigidity along the chain, which increases with chain length [25]. The angular distribution along the C-shape becomes increasingly uniform, thereby making the chain stiffer and more polarized along the x-axis. The RMS deviations from the the mean orientation thus decreases with system size (a larger fraction of the chain is polarized along the x-axis). This is in contrast to the positional steady-state fluctuations, where the mean $h(y)$ varies with chain length, and fluctuations about this exhibits roughness.
V.3 With finite rotational noise
Addition of rotational noise may indeed vary the possible DSS attained from simulation. We do not report them here; but simply mention that $\tau_{r}\approx 10$ sets a threshold above which the C-shape is the attractor. In this section we further investigate whether the aforementioned findings carry over to the case of finite $\tau_{r}$. Indeed we find that $\tau_{r}\gtrsim 20$ an identical FV scaling for $W_{h}$ is obtained, whereas $W_{\theta}$ only approximates the scaling law and collapse (15) - See Fig. 4. For $W_{h}$ one sees a crossover at $\tau_{r}$ for the angular fluctuations between a diffusive $\beta_{\theta}\approx 0.5$ and a deterministic $\beta_{\theta}\approx 1$. The effect of randomness prohibits a systems-spanning $l_{c}$. Nevertheless, a smoothness exponent with an identical scaling can be obtained.
For $\tau_{r}\lesssim 10$ there is no longer a propelling C-shape, and indeed other DSS are observed. The reason for this is that the C-shape DSS is sensitive to long-wavelength fluctuations in this regime, such that any fluctuations on the length scale $\frac{b^{2}D_{r}}{v_{s}}$ is greater than typical monomer lengths. It is to be noted that for equilibrated systems $\tau_{r}=\frac{8\pi\eta b^{3}}{k_{b}T}$, with $\tau_{t}=\frac{6\pi\eta b^{2}}{k_{b}T}$ ($\eta$ the viscosity of the medium). Thus the effect of rotational noise and hence potential loss of the C-shape DSS may set in for smaller $b$. This trade-off and its potential experimental relevance for the system considered here is discussed further in the Discussion.